OpenStudy (anonymous):

Half the difference of two numbers is 2. The sum of the greater number and twice the smaller number is 13. Find the numbers.

OpenStudy (lalaly):

let a=greater number and b=smaller Half the difference of two numbers is 2 so\[\frac{1}{2}(a-b)=2\] The sum of the greater number and twice the smaller number is 13 \[a+2b=13\]

OpenStudy (lalaly):

Do you know how to find the numbers now?

OpenStudy (anonymous):

I know, I know :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks

OpenStudy (lalaly):

hehe you are so smart fool.... and u ofcourse strawberry